# Geometric Framework for Cell Oversegmentation

### v2

@@ Appendix A Further Related Works
In this section, we provide further discussion of the 2D segmentation techniques used in CellPose and CellStitch, as well as some alternative 3D segmentation approaches.
2D Segmentation Methods. We introduce the following key 2D segmentation methods that facilitate the development of 3D cell segmentation methods:
@@
3D Segmentation Methods. We provide an overview of existing 3D segmentation techniques, along with their limitations in the context of the current state-of-the-art methods.
@@
Appendix B Technical Details
@@ B.1 Cross-layer Interpolation to Recover 2D Mis-segmented Cell Masks
B.2 Practical Framework to Deal with Tilted Oversegmentations
When tested with our end-users, we observed that handling tilted oversegmentation cases involved two stages of method transfer: (i) Cell type transfer, where the model transitions from in-domain plant cells to out-of-domain animal cells; and (ii) Geometric degradation, where some mask shape information is inevitably lost during the rotation and slicing process. To accommodate potential information loss, we propose a more flexible scheme that offers end-users a trade-off between precision and inclusiveness, allowing them to adjust the correction behavior based on their preference.
Penalties for Information Losses. As discussed, the reconstruction process may slightly distort the relative geometric positions of the cell masks, and interpolating between strips within each layer can alter the cross-layer overlapping areas and EMD values, thereby impacting both 3D shape indices and EMD-based features. To mitigate these information losses, we introduce a simple yet effective penalty mechanism applied to both geometric and topological features. Higher penalties suppress false positives by enforcing stricter consistency, while lower penalties allow greater flexibility, enabling the inclusion of borderline oversegmentation cases. These penalties can be tuned to match the specific objectives or tolerance levels of the end-user.
For the normalized 3D shape index retrieved from Algorithm 3, we apply a positive constant and use the penalized value as the input. For EMD information, recall that statistical features are extracted from each EMD array to ensure consistent input lengths. In our pre-trained model, we use the minimum and maximum as boundary features; here, the penalization can be adjusted by replacing the boundary features with quartiles ( and ). This approach implicitly penalizes the EMD information by narrowing the boundary features, making them stricter.
Experiment Results. We apply our framework to the 3D segmentation results generated by PlantSeg on the Pancreas-B dataset. As previously discussed, although as a 3D-based approach, PlantSeg performs relatively well on this dataset, likely due to the similar structure between cells in Pancreas-B and cells in its ovule-based training data. We conduct experiments under various penalty configurations to provide a reference for end-users. Nonetheless, we still strongly recommend that users incorporate human feedback to validate the final correction candidates, particularly if they opt to use our pre-trained model on plant cells without retraining it on their specific dataset.
| EMD Mode | Filtration | # Correct | # Unsure | # Incorrect | Accuracy | |
| 1.3 | 0.855 | 20 | 53 | 81 | 0.198 | |
| Min, Max | 2 | 0.871 | 19 | 48 | 71 | 0.211 |
| 2.5 | 0.882 | 18 | 44 | 64 | 0.220 | |
| 1.3 | 0.893 | 16 | 38 | 60 | 0.211 | |
| , | 2 | 0.920 | 14 | 25 | 47 | 0.230 |
| 2.5 | 0.941 | 11 | 17 | 35 | 0.240 |
As shown in Table 6, we adopt the human verification evaluation metric described in §4.2. Given the increased complexity as we need to handle all neighboring cells, we additionally report the filtration accuracy, which reflects the proportion of candidates excluded during screening. The penalty parameters serve as a trade-off between filtration rate (reducing human validation effort) and oversegmentation inclusion rate (capturing true positives). With appropriate penalization, only 5.9% of candidates are passed to human verification, offering end-users a flexible balance between efficiency and coverage.
B.3 Computational Efficiency
@@
MLP Hyperparameters. Our MLP is indeed lightweight and computationally efficient, which has 3 hidden layers with sizes 128, 64, and 32, uses ReLU activations, 0.3 dropout, a final sigmoid output.
B.4 Dealing with Multi-Oversegmentations
@@ C.1 More Successful Examples
Figure 8 illustrates an example of correcting a mis-segmentation that is not caused by an empty mask. In the CellStitch row, the purple cell at layer and the bright yellow cell at layer remain disconnected due to an undersegmentation error produced by CellPose2D at the same position in layer . Our method successfully identifies this oversegmentation from a 3D perspective and corrects the 2D undersegmentation. The raw image clearly indicates that the three cell masks belong to the same 3D cell body. Moreover, in line 6 of Algorithm 1, we mean that there is no “complete" cell existing between the gap, with both its highest and lowest layers located between our candidates. However, we allow noisy cell masks to exist between the candidates, enabling us to also address 2D undersegmentation errors.
@@
C.2 Inaccuracies in mAP and Jaccard Index
In Figure 11, we present two examples illustrating the limitations of mAP and Jaccard index, respectively, highlighting the necessity of human expert verification to validate the effectiveness of the corrections. The examples are drawn from CellPose-based results on plant cell datasets.
C.3 Examples for Unsure Case
@@
C.4 Analysis on Incorrect Case
@@
It is evident that both the dark blue and yellow masks only partially represent their corresponding ground truth cell masks. This highlights why we state that CellPose2D often “shrinks the cell masks smaller" than they are supposed to be. We illustrate this effect in Figure 15. When mask information is lost, both the geometric EMD measurement and the topological shape index are affected. For the EMD measurement, recall that EMD quantifies the effort required to transform one distribution into another. Losing geometric information alters the transformation, making the EMD between mask A and mask B differ from its original value. In the case shown in Figure 15, the EMD between the mis-segmented masks is smaller compared to the ground truth masks, suggesting a falsely higher similarity. For the topological shape index, which is calculated via changes in overlapping areas, as shown in Figure 14, the overlapping area between adjacent layers (e.g., layers and ) aligns more closely with the overlapping areas of preceding and succeeding layers (e.g., layers and ), as the masks shrink. These partial and altered representations lead our method to make incorrect judgments.
C.5 Comparison between PlantSeg and CellStitch on Pancreas-B Dataset
@@ Appendix D Dataset Information
Table 7 shows the information for both plant and animal dataset. All the plant-type datasets are publicly availiable at the following site [2]. For plant-type data, the voxel resolution is unknown. For animal-type data, the voxel resolution () is .
Pre-training Setup. For CellPose, we use the pre-trained cyto2 model to generate 2D pre-segmented results. We also tune the diameter case-by-case, from 60 to 100, based on the cells in each datasets. For PlantSeg, we use the pre-trained generic_confocal_3D_unet model to generate 3D segmentation results.
Appendix E Future Works and Limitations
Labeled Dataset for Pre-training. As discussed in previous sections, multi-stage transfer learning and geometric information loss can reduce the robustness of the correction framework. Constructing a pre-trained classifier tailored to the target dataset would significantly mitigate this issue. However, doing so typically requires substantial human-labeled data. To address this, future work could incorporate active learning strategies to automatically identify the most uncertain or influential oversegmentation cases for annotation. This approach would reduce labeling effort while enhancing generalization across diverse datasets.
Tackle Undersegmentations. Future work could also study how to extend the framework to also detect and correct undersegmentation (merging of multiple cells) and fragmentation (non-contiguous segments of a single cell), creating a more comprehensive post-processing tool with our work.
Broader Applications of Geo-Wasserstein Divergence. As a novel metric for capturing 2D geometric divergence, Geo-Wasserstein has shown promise in quantifying differences between cell masks. However, its potential remains underexplored. Future work could investigate its broader applications in other geometric domains, such as shape evolution, medical imaging, or morphometrics. Additionally, assessing its robustness across a wider range of cell types would be valuable. Due to the scarcity of publicly labeled 3D datasets, our evaluation was limited to plant and animal cells commonly used by end-users.

### v3

@@ Appendix A Further Related Works
In this section, we provide further discussion of the 2D segmentation techniques used in CellPose and CellStitch, as well as some alternative 3D segmentation approaches and their limitations when applying to our focused task.
2D Segmentation Methods.
We introduce the following key 2D segmentation methods that facilitate the development of 3D cell segmentation methods:
@@
3D Segmentation Methods.
We provide an overview of existing 3D segmentation techniques, along with their limitations in the context of the current state-of-the-art methods.
@@
Appendix B Technical Details for Implementation
@@ B.1 Cross-layer Interpolation to Recover 2D Mis-segmented Cell Masks
B.2 Computational Efficiency
@@
MLP Hyperparameters and Structure.
Our MLP is indeed lightweight and computationally efficient, which has 3 hidden layers with sizes 128, 64, and 32, uses ReLU activations, 0.3 dropout, a final sigmoid output.
B.3 Dealing with Multi-Oversegmentations
@@ C.1 More Successful Examples
Figure 8 illustrates an example of correcting a mis-segmentation that is not caused by an empty mask. In the CellStitch row, the purple cell at layer and the bright yellow cell at layer remain disconnected due to an undersegmentation error produced by CellPose2D at the same position in layer . Our method successfully identifies this oversegmentation from a 3D perspective and corrects the 2D undersegmentation. The raw image clearly indicates that the three cell masks belong to the same 3D cell body. Moreover, in line 6 of Algorithm 1, we mean that there is no “complete” cell existing between the gap, with both its highest and lowest layers located between our candidates. However, we allow noisy cell masks to exist between the candidates, enabling us to also address 2D undersegmentation errors.
@@
C.2 Examples for Unsure Case
@@
C.3 Analysis on Incorrect Case
@@
It is evident that both the dark blue and yellow masks only partially represent their corresponding ground truth cell masks. This highlights why we state that CellPose2D often “shrinks the cell masks smaller” than they are supposed to be. We illustrate this effect in Figure 13. When mask information is lost, both the geometric EMD measurement and the topological shape index are affected. For the EMD measurement, recall that EMD quantifies the effort required to transform one distribution into another. Losing geometric information alters the transformation, making the EMD between mask A and mask B differ from its original value. In the case shown in Figure 13, the EMD between the mis-segmented masks is smaller compared to the ground truth masks, suggesting a falsely higher similarity. For the topological shape index, which is calculated via changes in overlapping areas, as shown in Figure 14, the overlapping area between adjacent layers (e.g., layers and ) aligns more closely with the overlapping areas of preceding and succeeding layers (e.g., layers and ), as the masks shrink. These partial and altered representations lead our method to make incorrect judgments.
C.4 Comparison between PlantSeg and CellStitch on Pancreas-B Dataset
@@ Appendix D Dataset Information
Table 5 shows the information for both plant and animal dataset. All the plant-type datasets are publicly availiable at Bassel [2]. For plant-type data, the voxel resolution is unknown. For animal-type data, the voxel resolution () is .
Pre-training Setup.
For CellPose, we use the pre-trained cyto2 model to generate 2D pre-segmented results. We also tune the diameter case-by-case, from 60 to 100, based on the cells in each datasets. For PlantSeg, we use the pre-trained generic_confocal_3D_unet model to generate 3D segmentation results.